# Joint Pricing and Resource Allocation: An Optimal Online-Learning Approach

### v1

@@ 4.1 Algorithm Design Overview
We design a hierarchical algorithm to solve this problem. Algorithm 3 serves as the main algorithm, where we initialize $(mn+1)$ agents $\mathcal{A}_{K}$ for $K=0,1,\ldots,mn$ and run each of them for a period of time. The design of each agent $\mathcal{A}_{K}$ is defined as Algorithm 4, which contains the following three stages:
@@ 4.2 Meta-Algorithm: a Lower-Confidence-Bound (LCB) Strategy
Due to the piecewise convexity of $W(p)$, we divide the price range $[0,p_{\max}]$ into $(mn+1)$ intervals $[C_{i_{K},j_{K}},C_{i_{K+1},j_{K+1}}],K=0,1,\ldots,mn$. Within each interval, we initialize an OCO agent $\mathcal{A}_{K}$ that is responsible for converging to the local optimal. However, we cannot run multiple OCO agents simultaneously. Therefore, we require a meta-algorithm that serves as a manager over these agents and determine which $\mathcal{A}_{K}$ to run at each time, so as to locate the optimal price with the least cumulative regret.
@@
1:  Initialization: Set $\hat{p}_{K}^{*}:=C_{K,\tau},\hat{\vec{I}}^{*}_{K,0}=\vec{I}_{K,\tau,s}(C_{K,\tau})$ from Stage 1, and $\check{C}_{K}:=\log_{4/3}p_{\max}+1$.
2:  if $\Delta_{K}<+\infty$ then
3:     Define $N_{K,2}:=\frac{4\delta_{K}^{2}}{\Delta_{K}^{2}}$
4:  else
5:     Define $N_{K,2}:=n_{0}=6(\log_{4/3}T+\check{C}_{K})$
6:  end if
7:  Let $r_{K}:=\log_{2}(N_{K,2})$
8:  for $r=1,2,\ldots,r_{K}$ do
9:     for $t=1,2,\ldots,m_{r}:=2^{r}$ do
10:        Propose decisions $(\vec{I}_{t},p_{t})=(\hat{\vec{I}}_{K,r-1}^{*},\hat{p}_{K}^{*})$.
11:        Observe and record the marginal function $Q_{t}(\vec{I},\hat{p}_{K}^{*})$ with respect to $\vec{I}$.
12:     end for
13:     Set $\hat{\vec{I}}_{K,r}^{*}:=\mathop{\mathrm{argmin}}_{\vec{I}}\bar{Q}_{K,r}(\vec{I},\hat{p}_{K}^{*})$, where $\bar{Q}_{K,r}(\vec{I},\hat{p}_{K}^{*}):=\frac{1}{m_{r}}\cdot\sum_{t=1}^{\top}Q_{t}(\vec{I},\hat{p}_{K}^{*}).$
14:  end for
15:  Denote $\vec{I}_{K}^{*}:=\hat{\vec{I}}_{K,r_{K}}^{*},\hat{Q}_{K}^{*}:=\bar{Q}_{K,r_{K}}(\vec{I}_{K}^{*},\hat{p}_{K}^{*})$.
16:  Update $\hat{W}_{K}\leftarrow\hat{Q}_{K}^{*}-L_{W}\cdot\frac{1}{T}$ and $\Delta_{K}\leftarrow\frac{\delta_{K}}{\sqrt{N_{K,2}}}$.
@@
1:  Inherit $\hat{p}_{K}^{*},\hat{\vec{I}}_{K,r_{K}}^{*},\bar{Q}_{K,r_{K}}(\vec{I},\hat{p}_{K}^{*})$ from Algorithm 6.
2:  Denote $\vec{I}_{K}^{*}:=\hat{\vec{I}}_{K,r_{K}}^{*},\hat{Q}_{K}^{*}:=\bar{Q}_{K,r_{K}}(\vec{I}_{K}^{*},p_{K}^{*})$.
3:  Let $N_{K,3}\leftarrow N_{K,2}$ as its initialization.
4:  while $t\leq T$ do
5:     Propose decisions $(\vec{I}_{t},p^{\prime}_{t})=(\vec{I}^{*}_{K},\hat{p}_{K}^{*})$.
6:     Update $\hat{Q}_{K}^{*}\leftarrow\frac{N_{K,3}\cdot\hat{Q}_{K}^{*}+Q_{t}(\vec{I}_{t},p_{t})}{N_{K,3}+1}$, and $N_{K,3}\leftarrow N_{K,3}+1$
7:     Update $\hat{W}_{K}\leftarrow\hat{Q}_{K}^{*}-L_{W}\cdot\frac{1}{T}$ and $\Delta_{K}\leftarrow\frac{\delta_{K}}{\sqrt{N_{K,3}}}$.
8:  end while
@@ 4.3 Agent $\mathcal{A}_{K}$: a Zeroth-Order Optimizer
In this work, we propose Algorithm 4 as the framework of $\mathcal{A}_{K}$. In what follows, we elaborate each component of $\mathcal{A}_{K}$’s algorithmic design in detail.
Horizontal search space for $p_{K}^{*}$.
In the design of Stage $1$ algorithm as presented in Algorithm 5, we adopt the framework of zeroth-order online convex optimization. Specifically, we establish an epoch-based update rule of the search space of local optimal $p_{K}^{*}$. The search space (interval) for Epoch $\tau=1,2,\ldots$ is denoted as $[L_{K,\tau},U_{K,\tau}]$. Within each epoch, we divide the time horizon into a series of doubling sub-epochs to gather samples for $W(a),W(b),W(c)$ where $a,b,c$ are the three quarter points. By the end of each sub-epoch, we update the estimates and examine whether their estimation error bar is separable according to certain rules. As we keep doubling the size of sub-epochs, the estimation error bars are shrinking exponentially until they are separated. Then we reduce the search space by one quarter and proceed to Epoch $\tau+1$ Sub-Epoch $1$. When the search space is as sufficiently small as $O(1/T)$, we stop searching and proceed to Stage $2$.
Vertical uncertainty bound for $W(p_{K}^{*})$.
In Stage $1$, we maintain an error bar $\Delta_{K}$ as the confidence bound of estimating each local optimal $W(p_{K}^{*})$. We show that the error bar has a size of $\tilde{O}(\frac{1}{\sqrt{T_{K}}})$ if we have run $\mathcal{A}_{K}$ for $T_{K}$ times so far. In addition to the statistical concentrations, another intuition of this fact comes from Lemma A.9: If we cannot distinguish among the value of each quaternary of a convex function due to their uncertainty, then we may construct a comparable uncertainty bound for the optimal.
Complementary sampling to enhance $\Delta_{K}$.
It is worth noting that $\Delta_{K}<+\infty$ does not exist for granted even after Stage 1. This is because we cannot update $\Delta_{K}$ when the search space $[L_{K},U_{K}]$ is updated, i.e., no simultaneous “horizontal converging” and “vertical converging”. As a consequence, if we are very “lucky” that we can always reduce the search space in the first sub-epoch of every epoch until $U_{K,\tau}-L_{K,\tau}\leq 1/T$, then we will have $\Delta_{K}=+\infty$ until Stage $2$. We settle this issue in two approaches: (1) We upper bound the time periods before any $\Delta_{K}<+\infty$ by $O(\log T)$ based on the Pigeon-Hole Theorem. (2) We have Stage $2$ (Algorithm 6) as a complementary sampling stage without causing excessive regret. By the end of Stage $2$, we will have an ideal error bar for each agent $\mathcal{A}_{K}$.
Pure local exploitation contributing to global LCB.
From Agent $\mathcal{A}_{K}$’s perspective, it runs pure exploitation in Stage 3 (Algorithm 7) without causing extra sub-regret. However, it still keeps updating the estimates of $\hat{W}_{K}$ and $\Delta_{K}$ to facilitate the LCB meta-algorithm.
4.4 Technical Novelty
To the best of our knowledge, we are the first to propose a methodology undergoing “horizontal-and-vertical” convergence simultaneously. In contrast, existing works adopt either “vertical convergence” such as bandits algorithms (which allow non-convexity of objective functions but cannot achieve $O(\sqrt{T})$ regret even with smoothness assumptions), or “horizontal convergence” which is applicable to many online planning and optimization scenarios but requires global convexity assumptions.
@@ 5 Analysis
5.1 Sub-Regret and Confidence Bound of $\mathcal{A}_{K}$
In this subsection, we present two lemmas that show the convergence of each agent $\mathcal{A}_{K}$. Specifically, Lemma 5.2 shows the “horizontal convergence” of (inventory, price) decisions towards the local optimal. Lemma 5.3 shows the “vertical convergence” of estimation error $\Delta_{K}$ such that we are maintaining and updating a valid lower-confidence bound for the majority of time.
@@ 5.1 Sub-Regret and Confidence Bound of $\mathcal{A}_{K}$
For agent $\mathcal{A}_{K}$ defined as Algorithm 4 that has been running for $T_{K}$ time periods so far, the cumulative sub-regret is bounded by:
@@
The proof of Lemma 5.2 is relegated to Section A.5.
@@
The proof of Lemma 5.3 is presented in Section A.6. From Lemma 5.3, we directly get the following Corollary 5.4.
@@
The proof details of Lemma 5.5 is displayed in Section A.7.
5.2 Overall Regret
With the help of all the lemmas presented above, we are now ready to provide an upper bound on the total regret.

### v2

@@ 4.1 Algorithm Design Overview
We design a hierarchical algorithm to solve this problem. Algorithm 2 serves as the main algorithm. We firstly initialize $(mn+1)$ agents $\mathcal{A}_{K}$ for $K=0,1,\ldots,mn$ by running each of them for 3 times. After initialization, every time we select the $\hat{K}$-th agent $\mathcal{A}_{\hat{K}}$ whose Lower Confidence Bound $LCB_{\hat{K}}$ is the minimum, and run $\mathcal{A}_{\hat{K}}$ for a period of time (depending on the stage in which it is). The status of each $\mathcal{A}_{K}$ can be divided into three stages:
@@ 4.2 Meta-Algorithm: a Lower-Confidence-Bound (LCB) Strategy
Due to the piecewise convexity of $W(p)$, we divide the price range $[0,p_{\max}]$ into $(mn+1)$ intervals $[C_{i_{K},j_{K}},C_{i_{K+1},j_{K+1}}],K=0,1,\ldots,mn$. Within each interval, we initialize an OCO agent $\mathcal{A}_{K}$ that is responsible for converging to the local optimal. However, we cannot run multiple OCO agents simultaneously. Therefore, we require a meta-algorithm that serves as a manager over these agents and determine which $\mathcal{A}_{K}$ to run at each time, so as to locate the optimal price with the least regret.
@@
Algorithm 3 Agent $\mathcal{A}_{K}$ Stage 1
1:  Obtain $\hat{W}_{K},\Delta_{K}$ and £$\pounds_{K}$ from the Meta-Algorithm (Algorithm 2).
2:  for Epoch $\tau=1,2,\ldots,O(\log T)$do
3:     Let $a_{K,\tau}=\frac{3L_{K,1}+U_{K,1}}{4},c_{K,\tau}=\frac{L_{K,1}+U_{K,1}}{2},b_{K,\tau}=\frac{L_{K,1}+3U_{K,1}}{4}$.
4:     for  Sub-epoch $s=1,2,\ldots$ do
5:        Let sub-epoch length $n_{s}:=2^{s}$
6:        Define a $flag:=0$ for error-bar update.
7:        for $\hat{p}_{\tau}=a_{K,\tau},b_{K,\tau},c_{K,\tau}$ do
8:           for $t=1,2,\ldots,n_{s}$ do
9:              Propose decisions $(\vec{I}_{t},p_{t})=(\vec{I}_{K,\tau,s-1}(\hat{p}_{\tau}),\hat{p}_{\tau})$.
10:              Observe and record the marginal function $Q_{t}(\vec{I},\hat{p}_{\tau})$ with respect to $\vec{I}$.
11:           end for
12:           Define an aggregated function $Q_{K,\tau,s}(\vec{I},\hat{p}_{\tau}):=\frac{1}{n_{s}}\cdot\sum_{t=1}^{n_{s}}Q_{t}(\vec{I},\hat{p}_{\tau})$.
13:           Define the empirical optimal inventory $I_{K,\tau,s}(\hat{p}_{\tau}):=\mathop{\mathrm{argmin}}_{\vec{I}}Q_{K,\tau,s}(I,\hat{p}_{\tau})$.
14:           Denote $\hat{Q}_{K,\tau,s,\hat{p}_{\tau}}:=Q_{k,\tau,s}(I_{K,\tau,s}(\hat{p}_{\tau}),\hat{p}_{\tau})$, and $\Delta_{K,\tau,s}:=\frac{\delta_{K}}{2\sqrt{n_{s}}}$.
15:        end for
16:        if $\hat{Q}_{K,\tau,s,a_{K,\tau}}>\hat{Q}_{K,\tau,s,b_{K,\tau}}+4\Delta_{K,\tau,s}$ then
17:           Update $L_{K,\tau+1}\leftarrow a_{K,\tau}$, $U_{K,\tau+1}\leftarrow U_{K,\tau}$, $flag\leftarrow 1$.
18:        else if $\hat{Q}_{K,\tau,s,a_{K,\tau}}<\hat{Q}_{K,\tau,s,b_{K,\tau}}-4\Delta_{K,\tau,s}$ then
19:           Update $L_{K,\tau+1}\leftarrow L_{K,\tau}$, $U_{K,\tau+1}\leftarrow b_{K,\tau}$, $flag\leftarrow 1$.
20:        else if $\hat{Q}_{K,\tau,s,c_{K,\tau}}<\hat{Q}_{K,\tau,s,a_{K,\tau}}-4\Delta_{K,\tau,s}$ then
21:           Update $L_{K,\tau+1}\leftarrow a_{K,\tau}$, $U_{K,\tau+1}\leftarrow U_{K,\tau}$, $flag\leftarrow 1$.
22:        else if $\hat{Q}_{K,\tau,s,c_{K,\tau}}<\hat{Q}_{K,\tau,s,b_{K,\tau}}-4\Delta_{K,\tau,s}$ then
23:           Update $L_{K,\tau+1}\leftarrow L_{K,\tau}$, $b_{K,\tau+1}\leftarrow U_{K,\tau}$, $flag\leftarrow 1$.
24:        end if
25:        if $flag==1$ then
26:           if $U_{K,\tau+1}-L_{K,\tau+1}>\frac{1}{T}$ then
27:              Continue to Epoch $\tau+1$ (without updating $\hat{W}_{K}$ or $\Delta_{K}$).
28:           else
29:              Set $\hat{p}_{K}^{*}\leftarrow C_{K,\tau},\hat{\vec{I}}_{K,0}^{*}\leftarrow\vec{I}_{K,\tau,s}(C_{K,\tau})$.
30:              Update £$\pounds_{K}\leftarrow 2$ and Break (without updating $\hat{W}_{K}$ or $\Delta_{K}$).
31:           end if
32:        else if $\Delta_{K,\tau,s-1}<\Delta_{K}$ then
33:           Update $\Delta_{K}\leftarrow\Delta_{K,\tau,s-1},\hat{W}_{K}\leftarrow\min_{\hat{p}_{\tau}\in\{a_{K,\tau},c_{K,\tau},b_{K,\tau}\}}\hat{Q}_{K,\tau,s-1,\hat{p}_{\tau}}$.
34:        end if
35:     end for
36:  end for
@@
1:  Obtain $\hat{W}_{K},\Delta_{K}$ and £$\pounds_{K}$ from the Meta-Algorithm (Algorithm 2).
2:  Initialization: Set $\hat{p}_{K}^{*}:=C_{K,\tau},\hat{\vec{I}}^{*}_{K,0}=\vec{I}_{K,\tau,s}(C_{K,\tau})$ from Stage 1, and $\check{C}_{K}:=\log_{4/3}p_{\max}+1$.
3:  if $\Delta_{K}<+\infty$ then
4:     Let $N_{K,2}:=\frac{4\delta_{K}^{2}}{\Delta_{K}^{2}}$
5:  else
6:     Let $N_{K,2}:=n_{0}=6(\log_{4/3}T+\check{C}_{K})$
7:  end if
8:  Let $r_{K}:=\log_{2}(N_{K,2})$
9:  for $r=1,2,\ldots,r_{K}$ do
10:     for $t=1,2,\ldots,m_{r}:=2^{r}$ do
11:        Propose decisions $(\vec{I}_{t},p_{t})=(\hat{\vec{I}}_{K,r-1}^{*},\hat{p}_{K}^{*})$.
12:        Observe and record the marginal function $Q_{t}(\vec{I},\hat{p}_{K}^{*})$ with respect to $\vec{I}$.
13:     end for
14:     Set $\hat{\vec{I}}_{K,r}^{*}:=\mathop{\mathrm{argmin}}_{\vec{I}}\bar{Q}_{K,r}(\vec{I},\hat{p}_{K}^{*})$, where $\bar{Q}_{K,r}(\vec{I},\hat{p}_{K}^{*}):=\frac{1}{m_{r}}\cdot\sum_{t=1}^{\top}Q_{t}(\vec{I},\hat{p}_{K}^{*}).$
15:  end for
16:  Denote $\vec{I}_{K}^{*}:=\hat{\vec{I}}_{K,r_{K}}^{*},\hat{Q}_{K}^{*}:=\bar{Q}_{K,r_{K}}(\vec{I}_{K}^{*},\hat{p}_{K}^{*})$.
17:  Update $\hat{W}_{K}\leftarrow\hat{Q}_{K}^{*}-L_{W}\cdot\frac{1}{T},\Delta_{K}\leftarrow\frac{\delta_{K}}{\sqrt{N_{K,2}}}$ and £$\pounds_{K}\leftarrow 3$.
@@
1:  Obtain $\hat{W}_{K},\Delta_{K}$ and £$\pounds_{K}$ from the Meta-Algorithm (Algorithm 2).
2:  Initialization: Set $\hat{p}_{K}^{*},\hat{\vec{I}}_{K,r_{K}}^{*},\bar{Q}_{K,r_{K}}(\vec{I},\hat{p}_{K}^{*})$ from Stage 2 (Algorithm 4).
3:  Denote $\vec{I}_{K}^{*}:=\hat{\vec{I}}_{K,r_{K}}^{*},\hat{Q}_{K}^{*}:=\bar{Q}_{K,r_{K}}(\vec{I}_{K}^{*},p_{K}^{*})$.
4:  Let $N_{K,3}\leftarrow N_{K,2}$ as its initialization.
5:  while $t\leq T$ do
6:     Propose decisions $(\vec{I}_{t},p^{\prime}_{t})=(\vec{I}^{*}_{K},\hat{p}_{K}^{*})$.
7:     Update $\hat{Q}_{K}^{*}\leftarrow\frac{N_{K,3}\cdot\hat{Q}_{K}^{*}+Q_{t}(\vec{I}_{t},p_{t})}{N_{K,3}+1}$, and $N_{K,3}\leftarrow N_{K,3}+1$
8:     Update $\hat{W}_{K}\leftarrow\hat{Q}_{K}^{*}-L_{W}\cdot\frac{1}{T}$ and $\Delta_{K}\leftarrow\frac{\delta_{K}}{\sqrt{N_{K,3}}}$.
9:  end while
@@ 4.3 Agent $\mathcal{A}_{K}$: a Zeroth-Order Optimizer
We present a detailed introduction of each component of $\mathcal{A}_{K}$ in Appendix A.
Technical Novelty We propose a unique methodology undergoing “horizontal-and-vertical” convergence simultaneously, for the first time. In contrast, existing works adopt either “vertical convergence” such as bandits algorithms (which allow non-convexity of objective functions but cannot achieve $O(\sqrt{T})$ regret even with smoothness assumptions), or “horizontal convergence” which is applicable to many online planning and optimization scenarios but requires global convexity assumptions.
@@ 5 Analysis
In the following, we present two lemmas that show the convergence of each agent $\mathcal{A}_{K}$. Specifically, Lemma 5.2 shows the “horizontal convergence” of (inventory, price) decisions towards the local optimal. Lemma 5.3 shows the “vertical convergence” of estimation error $\Delta_{K}$ such that we are maintaining and updating a valid lower-confidence bound for the majority of time.
@@
For agent $\mathcal{A}_{K}$ (defined as Algorithms 3, 4 and 5) that has been running for $T_{K}$ time periods so far, the cumulative sub-regret is bounded by:
@@
The proof of Lemma 5.2 is relegated to Section B.4.
@@
The proof of Lemma 5.3 is in Section B.5. From Lemma 5.3, we directly get the following corollary.
@@
The proof details of Lemma 5.5 is displayed in Section B.6. With the help of all the lemmas above, we are now ready to provide an upper bound on the total regret.